# Elastic bounds of the coupling tensor of anisotropic composite laminates: a polar approach

### v2

@@ 1 Introduction
$\left\{\begin{array}[]{c}\mathbf{N}\\ \mathbf{M}\end{array}\right\}=\left[\begin{array}[]{cc}h\mathbb{A}&\dfrac{h^{2}}{2}\mathbb{B}\\ \dfrac{h^{2}}{2}\mathbb{B}&\dfrac{h^{3}}{12}\mathbb{D}\end{array}\right]\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\},$ (1)
@@ 3 Statement of the problem
$U=\frac{1}{2}\left\{\begin{array}[]{c}\textbf{N}\\ \textbf{M}\end{array}\right\}\cdot\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\}=\frac{1}{2}\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\}\cdot\left[\begin{array}[]{cc}h\mathbb{A}&\dfrac{h^{2}}{2}\mathbb{B}\\ \dfrac{h^{2}}{2}\mathbb{B}&\dfrac{h^{3}}{12}\mathbb{D}\end{array}\right]\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\},$ (5)
@@ 5.1 Aligned orthotropic tensors
Because $T_{1}>0$ and by eqs. (43)${}_{1}$ and (44)${}_{2}$, condition (43)${}_{2}$ is redundant and the above conditions can be rewritten as
@@
A scrutiny of $M4_{2}$ and of $M4_{3}$ allows to see that if conditions (45)${}_{1,3}$ are considered in eqs. (42)${}_{2,3}$, this will lead to two conditions that are exactly the analogous of those that could be obtained by eqs.(20) and (32), i.e. writing the conditions $M2$ and $M3$ also for $\mathbb{D}$.
@@ 5.2 Square symmetric $\mathbb{B}$
Let us now consider the case of a square-symmetric coupling $\mathbb{B}$, i.e. a laminate designed to have $R_{1}^{B}=0$. In such a case, eq. (36)${}_{9}$ reduces to
@@
which is redundant, as previously discussed in Sect. 4.1. Moreover, eq. (36)${}_{10}$ becomes exactly eq. (36)${}_{8}$, so it can be discarded too. Actually, in this special case it is the whole set of conditions on $M3$ that it is redundant. This is quite natural, because when $R_{1}^{B}=0$ the only parameter accounting for coupling disappears from $M3$, which reduces by consequence to conditions already accounted for by $M2$. Moreover, eq. (35) reduces to
@@ 5.4 $R_{0}$-orthotropic laminates
If once again we consider the case of aligned tensors, eq. (38), then proceeding like in Sect. 5.1, we get that condition (55)${}_{4}$ is redundant and finally
@@ 5.5 Coupled isotropic laminates
eq. (36)${}_{9}$ is redundant with respect to eq. (36)${}_{10}$ and finally conditions (36) reduce to only

### v3

@@ 1 Introduction
$\left\{\begin{array}[]{c}\mathbf{N}\\ \mathbf{M}\end{array}\right\}=\left[\begin{array}[]{cc}h\mathbb{A}&\dfrac{h^{2}}{2}\mathbb{B}\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \dfrac{h^{2}}{2}\mathbb{B}&\dfrac{h^{3}}{12}\mathbb{D}\end{array}\right]\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\},$ (1)
@@ 3 Statement of the problem
$U=\frac{1}{2}\left\{\begin{array}[]{c}\textbf{N}\\ \textbf{M}\end{array}\right\}\cdot\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\}=\frac{1}{2}\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\}\cdot\left[\begin{array}[]{cc}h\mathbb{A}&\dfrac{h^{2}}{2}\mathbb{B}\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \dfrac{h^{2}}{2}\mathbb{B}&\dfrac{h^{3}}{12}\mathbb{D}\end{array}\right]\left\{\begin{array}[]{c}\bm{\varepsilon}\\ \bm{\kappa}\end{array}\right\},$ (5)
@@ 5.1 Aligned orthotropic tensors
Because $T_{1}>0$ and by eqs. (43)1 and (44)2, condition (43)2 is redundant and the above conditions can be rewritten as
@@
A scrutiny of $M4_{2}$ and of $M4_{3}$ allows to see that if conditions (45)1,3 are considered in eqs. (42)2,3, this will lead to two conditions that are exactly the analogous of those that could be obtained by eqs.(20) and (32), i.e. writing the conditions $M2$ and $M3$ also for $\mathbb{D}$.
@@ 5.2 Square symmetric $\mathbb{B}$
Let us now consider the case of a square-symmetric coupling $\mathbb{B}$, i.e. a laminate designed to have $R_{1}^{B}=0$. In such a case, eq. (36)9 reduces to
@@
which is redundant, as previously discussed in Sect. 4.1. Moreover, eq. (36)10 becomes exactly eq. (36)8, so it can be discarded too. Actually, in this special case it is the whole set of conditions on $M3$ that it is redundant. This is quite natural, because when $R_{1}^{B}=0$ the only parameter accounting for coupling disappears from $M3$, which reduces by consequence to conditions already accounted for by $M2$. Moreover, eq. (35) reduces to
@@ 5.4 $R_{0}$-orthotropic laminates
If once again we consider the case of aligned tensors, eq. (38), then proceeding like in Sect. 5.1, we get that condition (55)4 is redundant and finally
@@ 5.5 Coupled isotropic laminates
eq. (36)9 is redundant with respect to eq. (36)10 and finally conditions (36) reduce to only